# Optimizing the image projection of spatially incoherent light from a multimode fiber

### v1

@@ I Introduction
Though lasers with both large temporal and spatial coherence continue to be a hot topic in various fields, light source with a reduced spatial coherence has also been gaining attentions these years. Spatially incoherent light has the natural property of resisting the external perturbations, and has been demonstrated potential applications in many fields, for example in free-space optical communicationsShirai et al. (2003); Gbur (2014) and light potential projections for ultracold atomsWei (2019).
@@
To reduce the spatial incoherence, one conventional approach is based on the time averaging, including both active and passive methodsWei (2019). However, at any particular time point, the light field still maintains long-distance correlation. Thus the claim of an effective spatially incoherent light actually depends on the typical time scale of the targeted system. In other words, if the typical evolution frequency of the targeted system is faster than the time averaging frequency in producing spatial incoherence, this method will become invalid.
@@
The targeted system that we are interested in is the ultracold atom gas, loaded into an optical dipole trap or an optical lattice Schafer et al. (2020); Gross and Bloch (2017). We intend to project arbitrary light potentials generated by a digital micromirror device (DMD) onto atoms to realize manipulations of quantum states Di Carli et al. (2024); Amico et al. (2021, 2022); Zupancic et al. (2016); Tajik et al. (2019). In practice, unavoidable dusts, roughness, or reflections in the optical path will significantly distort the projected shapes and cause unwanted interference patterns if we use a coherent laser. In the past decade, this issue has been shown to be possibly solved by using a spatially incoherent light to illuminate the DMD chip Braun et al. (2024); Hirthe (2022) instead.
@@
The time scale of our interested system is determined by the optical potentials depth, typically about several $10$ kHz Bakr (2011) for optical lattices. Thus the typical frequency in generating the spatial incoherence must be high. Additionally, the ultracold atom gas in an optical lattice is sensitive to noise and perturbations, so stability is also concerned.
@@
During the past decades, various ways of generating a spatial incoherent light based on the time averaging idea have been proposed and developed. A diffuser rotated by a motor is maybe the simplest way to destroy the spatial coherence Stangner et al. (2017). And people can also use a galvo or an acousto optical deflector (AOD) to modulate the incident angle shining on a static diffuser to do so Wei (2019). Another way is to use a multimode fiber (MMF) to turn the temporal incoherence into the spatial incoherenceSaleh and Teich (Wiley, 2019), which is the main focus of this work. Actually, the MMF can also be combined with other techniques (including diffusers, galvos, AODs, vibrators, and etc Xu et al. (2023); Mehta et al. (2012); Mao et al. (2023); Manni and Goodman (2012); Tong et al. (2017); Ma et al. (2016); Ha et al. (2009); Yang et al. (2024)) to be applied in many situations. But for our specific purpose, we only discuss about the scenario of a static MMF associated with a LED source, for its several advantages:
@@
Properties of the MMF in the process of generating the spatial incoherence has already been studied these years, both experimentally and theoreticallyDandliker et al. (1985); Efimov (2014, 2019a, 2019b, 2019c, 2018). We extent the former models in the context of DMD projection, trying to bridge between the previous extensive studies and the applications in the ultracold-atom experiments. A full-numerical model of the MMF, which is also compatible with other optical process including the free space propagation, is established, and part of the results is compared with our experimental data. We mainly investigate the intrinsic speckle contrast and the spatial correlation functions (or the perturbation robustness) of the MMFs in the projection process, which will be discussed in detail in the following sections.
@@ II MMF projection settings and the numerical model
To closely study the properties of the light injected onto the DMD for use, we establish a full-numerical model to simulate the light propagation in the MMF, as shown in Figure 2. The intensity distribution at the MMF output (fiber endface) is the result of interference between all the radial modes, whose phases are not fixed with each other due to the fiber time delay and the finite temporal coherence length. An example of the fiber-supported radial modes are shown in Figure 2(a), by numerically solving the Helmholtz equationSaleh and Teich (Wiley, 2019):
@@
The complex amplitude function $U(\vec{r})$ can be decomposed into radial, angular, and axial components as: $U=u_{m}(r)e^{-jl\phi}e^{-j\beta_{lm}z}$, where $r$ and $\phi$ are radial coordinates in the fiber cross section, and $z$ is in the light propagation (axial) direction. $l=0,~{}\pm 1,~{}\pm 2,...$ represents the angular momentum, and $\beta_{lm}$ is the axial wave-number. We use $U_{lm}$ to label the radial mode with the angular momentum $l$ and the axial wave-number $\beta_{lm}$ as in Saleh and Teich (Wiley, 2019). In the following, we mainly use a polarized model in analysis, where the polarization degeneracy of each $U_{lm}$ is neglectedEfimov (2019c), while a vector version of our model is also taken into comparison in Figure5. The total number of the radial modes supported by a MMF is determined by its $NA$ and the size of core, and we have more than $2200$ modes for a standard MMF with $NA=0.22$ and core $D=105~{}\mu$m.
@@
In the following, we calculate the intensity distributions at the endface for different fiber parameters. There are two observables extracted from the intensity distribution that can reflect the overall properties of the MMF: the intrinsic speckle contrast and the spatial correlation function Efimov (2014, 2019a, 2019b, 2019c).
@@
Our model also simulates the light propagation in the free space which is included in the common DMD projection setting. The light propagation is governed by the Rayleigh-Sommerfeld integral formulaErsoy (Wiley, 2006):
@@
We note that, compared to existing modelsEfimov (2019c, 2014), we provide here a full numerical method to investigate various properties of the partially temporal-spatially incoherent light generated from a MMF conveniently and directly. This model enables us not only to look into properties on the fiber endface, but also to study the propagation after the MMF, especially in the DMD projection setting.
@@ III Results: Numerical and experimental
MMF length and temporal coherence. Our numerical model enables us to study the influence of both the fiber length and the temporal coherence length of the light source. Actually, the parameter really matters here is the ratio of the total time delay $t$ caused by the MMF to the temporal coherence length $\tau_{c}$. The MMF caused total time delay is expressed asSaleh and Teich (Wiley, 2019):
@@
Figure 3 shows the dependence of both the intrinsic speckle contrast and the spatial correlation function on this dimensionless parameter $t/\tau_{c}$. In Figure 3, the MMF parameters (NA, core radius, and fiber length) are fixed and we only change the temporal coherence $\tau_{c}$ of the light source (see Figure 2(b)). Figure 3 reveals that both the intrinsic speckle contrast and the residual coherence decrease monotonously with $t/\tau_{c}$, as expected intuitively. The corresponding spatially correlation functions and endface intensity distributions are demonstrated in Figure 3(b) and (c). Results above are in good agreement with former modelsEfimov (2019c). We note here that, the saturation behavior at the large $t/\tau_{c}$ side is a numerical effect, and the saturation limit is limited by the finite sampling time in our model, as shown in the inset of Figure 3(a).
@@
MMF radius and robustness against perturbation. We now study how the performance of a MMF is influenced by its core radius. A MMF with a larger core radius will support more radial modes in the scale of $R^{2}$aleh and Teich (Wiley, 2019). In the meantime, the total time delay through the MMF is only determined by its NA, which is fixed at $0.22$ in our discussion. Thus a larger core radius results in a denser mode distribution along the time axis. Intuitively, this will influence the dependence of the intrinsic speckle contrast on the parameter $t/\tau_{c}$ (see Figure 3), resulting in a later saturation. On the other hand, the existence of more radial modes means that the superposition of the light field will become more chaotic, and may lead to a better robustness against perturbations in the real optical paths. In the following, we compare the MMFs with various core radii in these two aspects.
@@
A further question arises here that, whether the residual coherence is the only important and sufficient indicator to summarize the spatial correlation function of a MMF? In Figure 7, we point out that on the image plane of the magnifying 4f-system (where we plane to place the DMD chip), the spatial coherence length Wei (2019):

### v2

@@ I Introduction
Though lasers with both large temporal and spatial coherence continue to be a hot topic in various fields, light source with a reduced spatial coherence has also been gaining attentions these years. Spatially incoherent light has the natural property of resisting the external perturbations, and has been demonstrated potential applications in many fields, for example in free-space optical communications[1, 2] and light potential projections for ultracold atoms[3].
@@
To reduce the spatial incoherence, one conventional approach is based on the time averaging, including both active and passive methods[3]. However, at any particular time point, the light field still maintains long-distance correlation. Thus the claim of an effective spatially incoherent light actually depends on the typical time scale of the targeted system. In other words, if the typical evolution frequency of the targeted system is faster than the time averaging frequency in producing spatial incoherence, this method will become invalid.
@@
The targeted system that we are interested in is the ultracold atom gas, loaded into an optical dipole trap or an optical lattice [4, 5]. We intend to project arbitrary light potentials generated by a digital micromirror device (DMD) onto atoms to realize manipulations of quantum states [6, 7, 8, 9, 10]. In practice, unavoidable dusts, roughness, or reflections in the optical path will significantly distort the projected shapes and cause unwanted interference patterns if we use a coherent laser. In the past decade, this issue has been shown to be possibly solved by using a spatially incoherent light to illuminate the DMD chip [11, 12] instead.
@@
The time scale of our interested system is determined by the optical potentials depth, typically about several $10$ kHz [13] for optical lattices. Thus the typical frequency in generating the spatial incoherence must be high. Additionally, the ultracold atom gas in an optical lattice is sensitive to noise and perturbations, so stability is also concerned.
@@
During the past decades, various ways of generating a spatial incoherent light based on the time averaging idea have been proposed and developed. A diffuser rotated by a motor is maybe the simplest way to destroy the spatial coherence [14]. And people can also use a galvo or an acousto optical deflector (AOD) to modulate the incident angle shining on a static diffuser to do so [3]. Another way is to use a multimode fiber (MMF) to turn the temporal incoherence into the spatial incoherence[15], which is the main focus of this work. Actually, the MMF can also be combined with other techniques (including diffusers, galvos, AODs, vibrators, and etc [16, 17, 18, 19, 20, 21, 22, 23]) to be applied in many situations. But for our specific purpose, we only discuss about the scenario of a static MMF associated with a LED source, for its several advantages:
@@
Properties of the MMF in the process of generating the spatial incoherence has already been studied these years, both experimentally and theoretically[24, 25, 26, 27, 28, 29]. We extent the former models in the context of DMD projection, trying to bridge between the previous extensive studies and the applications in the ultracold-atom experiments. A full-numerical model of the MMF, which is also compatible with other optical process including the free space propagation, is established, and part of the results is compared with our experimental data. We mainly investigate the intrinsic speckle contrast and the spatial correlation functions (or the perturbation robustness) of the MMFs in the projection process, which will be discussed in detail in the following sections.
@@ II MMF projection settings and the numerical model
To closely study the properties of the light injected onto the DMD for use, we establish a full-numerical model to simulate the light propagation in the MMF, as shown in Figure 2. The intensity distribution at the MMF output (fiber endface) is the result of interference between all the radial modes, whose phases are not fixed with each other due to the fiber time delay and the finite temporal coherence length. An example of the fiber-supported radial modes are shown in Figure 2(a), by numerically solving the Helmholtz equation[15]:
@@
The complex amplitude function $U(\vec{r})$ can be decomposed into radial, angular, and axial components as: $U=u_{m}(r)e^{-jl\phi}e^{-j\beta_{lm}z}$, where $r$ and $\phi$ are radial coordinates in the fiber cross section, and $z$ is in the light propagation (axial) direction. $l=0,~{}\pm 1,~{}\pm 2,...$ represents the angular momentum, and $\beta_{lm}$ is the axial wave-number. We use $U_{lm}$ to label the radial mode with the angular momentum $l$ and the axial wave-number $\beta_{lm}$ as in [15]. In the following, we mainly use a polarized model in analysis, where the polarization degeneracy of each $U_{lm}$ is neglected[28], while a vector version of our model is also taken into comparison in Figure5. The total number of the radial modes supported by a MMF is determined by its $NA$ and the size of core, and we have more than $2200$ modes for a standard MMF with $NA=0.22$ and core $D=105~{}\mu$m.
@@
In the following, we calculate the intensity distributions at the endface for different fiber parameters. There are two observables extracted from the intensity distribution that can reflect the overall properties of the MMF: the intrinsic speckle contrast and the spatial correlation function [25, 26, 27, 28].
@@
Our model also simulates the light propagation in the free space which is included in the common DMD projection setting. The light propagation is governed by the Rayleigh-Sommerfeld integral formula[30]:
@@
We note that, compared to existing models[28, 25], we provide here a full numerical method to investigate various properties of the partially temporal-spatially incoherent light generated from a MMF conveniently and directly. This model enables us not only to look into properties on the fiber endface, but also to study the propagation after the MMF, especially in the DMD projection setting.
@@ III Results: Numerical and experimental
MMF length and temporal coherence. Our numerical model enables us to study the influence of both the fiber length and the temporal coherence length of the light source. Actually, the parameter really matters here is the ratio of the total time delay $t$ caused by the MMF to the temporal coherence length $\tau_{c}$. The MMF caused total time delay is expressed as[15]:
@@
Figure 3 shows the dependence of both the intrinsic speckle contrast and the spatial correlation function on this dimensionless parameter $t/\tau_{c}$. In Figure 3, the MMF parameters (NA, core radius, and fiber length) are fixed and we only change the temporal coherence $\tau_{c}$ of the light source (see Figure 2(b)). Figure 3 reveals that both the intrinsic speckle contrast and the residual coherence decrease monotonously with $t/\tau_{c}$, as expected intuitively. The corresponding spatially correlation functions and endface intensity distributions are demonstrated in Figure 3(b) and (c). Results above are in good agreement with former models[28]. We note here that, the saturation behavior at the large $t/\tau_{c}$ side is a numerical effect, and the saturation limit is limited by the finite sampling time in our model, as shown in the inset of Figure 3(a).
@@
MMF radius and robustness against perturbation. We now study how the performance of a MMF is influenced by its core radius. A MMF with a larger core radius will support more radial modes in the scale of $R^{2}$[15]. In the meantime, the total time delay through the MMF is only determined by its NA, which is fixed at $0.22$ in our discussion. Thus a larger core radius results in a denser mode distribution along the time axis. Intuitively, this will influence the dependence of the intrinsic speckle contrast on the parameter $t/\tau_{c}$ (see Figure 3), resulting in a later saturation. On the other hand, the existence of more radial modes means that the superposition of the light field will become more chaotic, and may lead to a better robustness against perturbations in the real optical paths. In the following, we compare the MMFs with various core radii in these two aspects.
@@
A further question arises here that, whether the residual coherence is the only important and sufficient indicator to summarize the spatial correlation function of a MMF? In Figure 7, we point out that on the image plane of the magnifying 4f-system (where we plane to place the DMD chip), the spatial coherence length [3]: